# Shape-induced Bose-Einstein Condensation

### v1

@@ Abstract
Recent advancements in flat-bottomed optical box traps have enabled the realization of homogeneous Bose gases, allowing for the exploration of Bose-Einstein condensation in more complex confinement geometries. Here we propose a shape-induced Bose-Einstein condensation at fixed size, temperature, and density. We investigate the impact of quantum shape effects on a non-interacting Bose gas confined within nested square domains, where the shape parameter is defined and controlled by the rotation angle between the inner and outer squares. The size-invariant shape transformation significantly affects the condensate fraction and specific heat of the system. Our work opens new avenues for controlling quantum systems through geometric manipulation, providing insights into the thermodynamic properties of Bose gases under shape-induced quantum effects.
@@
Quantum-confined particles in box-like potentials with well-defined size parameters are often discussed in the context of quantum size effects, where energy quantization becomes prominent as the thermal de Broglie wavelength of particles, $\lambda_{th}=\hbar\sqrt{2\pi}/\sqrt{mk_{B}T}$, approaches the box size [15, 16]. The size parameters are defined under Lebesgue measure as volume, surface area, periphery, and the number of vertices [17, 18], which are collectively called Weyl size parameters. Both the dimensionality and sizes are critical control parameters for the thermodynamic properties of a Bose gas as well as the BEC transition. Quantum size effects in Bose-Einstein condensates have been explored extensively in harmonic traps [19, 20, 21, 22, 23, 24, 25, 26, 27], rotating traps [28], systems with finite number of particles [29, 30, 31], low-dimensional traps[32, 33, 34], and anisometric traps with multistep condensations [35, 36, 37, 38]. However, a more intriguing scenario emerges when the size of the confining potential remains constant, but its shape is altered at fixed temperature and density.
@@
While previous studies have investigated BEC in various arbitrarily shaped potentials [39, 40, 41, 42], these cannot be considered as pure shape effects, because they do not ensure size-invariance; inevitably, one or more size parameters change, making the size effects unavoidably vary. Conversely, a novel quantum-mechanical phenomenon known as the quantum shape effect occurs when the geometry of the box potential is modified while keeping all the Lebesgue size parameters, as well as the topology and boundary curvature of the confinement domain, unchanged [43, 44, 45]. The quantum shape effect is achieved through a geometric technique called size-invariant shape transformation [46], illustrated in Fig. 1.
@@
In this Letter, we investigate the condensation of a non-interacting Bose gas with a finite number of particles confined in a 2D flat-bottom potential between nested square domains with impenetrable boundaries. The angle between the sides of the inner and outer squares defines a size-invariant shape parameter. We demonstrate that quasistatic changes in a shape parameter can induce the BEC. The condensate fraction can be changed and controlled just by shape, without altering the size parameters of the confinement domain, as well as the density and temperature of the gas. Then, we investigate the behavior of Bose-Einstein gas under different dimensionless temperatures and densities to find the optimum condition where the change in condensate fraction due to shape effects is maximized. We calculate the condensate fraction and heat capacity of a Bose gas changing with shape, temperature, and density. Note that our focus is solely on demonstrating that shape alone can induce the condensation of a Bose gas. As long as the overall confinement is sufficiently strong, the effects of geometry can be made dominant over inter-particle interactions [49, 50, 51, 44]. Consequently, we do not employ the Gross-Pitaevskii equation, as the dynamics or other properties of the condensate are beyond the scope of this study.
@@
Confinement geometry.—A non-interacting Bose gas is confined in the region between two nested square domains as shown in Fig. 1 at $\tilde{T}=1$ and $\tilde{n}_{r}=1$, with the potential being infinite inside the inner square and outside the outer square, and zero in between. The outer boundaries of the domain are defined by the side length $L_{o}=L$, while the inner boundaries are set to $L_{i}=0.7L$. The shape control is governed by the parameter $\theta$, which represents the rotation angle of the inner core, allowing us to make size-invariant changes to the domain. Since our investigation focuses explicitly on the BEC transition, we employ a quantum statistical mechanical framework. The confined Bose gas obeys the Bose-Einstein distribution, $f=1/\{\exp[(E-\mu)/(k_{B}T)]-1\}$. There is no explicit time variable, and all changes to the system are performed quasistatically to ensure that thermal equilibrium is maintained at all times. Hence, to determine the energy eigenvalues, we numerically solved the time-independent Schrödinger equation, $\hat{H}\Psi=E\Psi$, for the domain $\Omega$. The Hamiltonian of the system includes explicit dependence on the shape parameter, $\theta$,
@@
To make our results independent of the specific type of Bose gas and any other free parameters, we employed dimensionless values throughout this work. Initially, we selected Helium-4 mass and $L=100$ nm for our calculations. Using these initial values, we calculated the discrete energy eigenvalues for each degree step from $\theta=0$ to $\theta=45$ degrees. We use 5000 of them for each case to ensure the partition function saturates due to the soft cutoff provided by the Bose function. By doing so, we obtained the energy spectrum for each angular configuration. Subsequently, we converted these results into dimensionless values to generalize our findings and eliminate dependence on the chosen initial parameters.
The dimensionless temperature is defined as $\tilde{T}=T/T_{0}$, where $T_{0}$ is the reference temperature obtained from the ground state energy $E_{0}$ at $\theta=45^{\circ}$ configuration (i.e. minimum $E_{0}$ value), $T_{0}=E_{0}/k_{B}$. Energies are normalized by $k_{B}T$ as, $\tilde{E}=E/(k_{B}T)$, and the dimensionless density is defined as $\tilde{n}_{r}=n/n_{q}(T_{0})=\lambda_{th}^{2}(T_{0})N/A$, where $n=N/A$ is classical surface density, $N$ is the number of particles and $A$ is the area of the confinement domain $\Omega$. Here, $n_{q}(T)=1/\lambda_{th}^{2}(T)$ is the quantum density. The dimensionless density is renormalized using the quantum density at the reference temperature to ensure that the density, and thus the number of particles remain constant for different temperatures.
Condensate fraction.—Fig. 1 provides qualitative support for the appearance of condensation with increasing $\theta$, indicated by the increased probability density of the ground state. To analyze the BEC transition due to shape effect quantitatively, we determine the condensate fraction,
@@
where $N_{GS}$ is the number of particles in the ground state. $\sum_{E_{GS}}$ represents the sum over the (fourfold) degenerate ground state energy levels, whereas $\sum_{E}$ represents the sums over all energy levels. The chemical potential is calculated numerically as the inverse solution for a given particle density and temperature.
@@
Specific heat.—Next, we examine the dimensionless specific heat capacity at constant Weyl size parameters (volume, area, periphery, vertices), denoted by $W$ subscript in Eq. (3), under the change of $\theta$ for dimensionless density $\tilde{n}_{r}=1$ and dimensionless temperatures $\tilde{T}=\{0.1,0.5,1,5\}$ as seen in Fig. 3a. The specific heat exhibits distinct and intriguing characteristics that differ significantly from those observed in Bose gases confined in regular flat-bottom geometries or harmonic traps, even in its temperature variation, Fig. 3b.
@@
Conclusion.—In conclusion, we have introduced the phenomenon of shape-induced BEC. Size-invariant shape transformations lead to unique behaviors in the condensate fraction and specific heat, offering additional control over the properties of Bose-Einstein condensates. Our findings reveal that the condensate fraction is significantly influenced by the shape parameter. Moreover, we have observed that the shape-dependent specific heat exhibits markedly different behaviors compared to conventional non-interacting Bose gases confined in other potentials.

### v2

# Shape-controlled Bose-Einstein Condensation

@@ Abstract
Size-invariant shape transformation is a geometric technique that allows for a clear separation between quantum size and shape effects by modifying the shape of the confinement domain without altering its size. The impact of shape on the behavior of confined systems is significantly different from that of size, making it an emerging area of research. The recent realization of flat-bottomed optical box traps has further contributed to the study of quantum gases in complex confinement geometries. Here, we propose shape-induced Bose-Einstein condensation at a fixed size, temperature, and density. We investigate the impact of pure quantum shape effects on a non-interacting Bose gas confined within nested square domains, where the shape parameter is defined and controlled by the rotation angle between the inner and outer squares. Our findings reveal that specific heat exhibits an additional low-temperature peak at certain shapes. This work opens new avenues for controlling quantum systems through geometric manipulation and provides insights into the thermodynamic properties of Bose gases under shape-induced quantum effects.
I Introduction
@@
Quantum-confined particles in box-like potentials with well-defined size parameters are often discussed in the context of quantum size effects, where energy quantization becomes prominent as the thermal de Broglie wavelength of particles, $\lambda_{th}=\hbar\sqrt{2\pi}/\sqrt{mk_{B}T}$, approaches the box size [15, 16]. The size parameters are defined under the Lebesgue measure as volume, surface area, periphery, and the number of vertices [17, 18], which are collectively called Weyl size parameters. Both the dimensionality and sizes are critical control parameters for the thermodynamic properties of a Bose gas as well as the BEC transition. Quantum size effects in Bose-Einstein condensates have been explored extensively in harmonic traps [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], rotating traps [30], systems with finite number of particles [31, 32, 33, 34, 35], low-dimensional traps[36, 37, 38, 39], and anisometric traps with multistep condensations [40, 41, 42, 43]. However, a more intriguing scenario emerges when the size of the confining potential remains constant, but its shape is altered at fixed temperature and density.
@@
While previous studies have investigated BEC in various arbitrarily shaped potentials [44, 45, 46, 47, 48, 49, 50], these cannot be considered as pure shape effects, because they do not ensure size-invariance; inevitably, one or more size parameters change, making the size effects unavoidably vary. Conversely, a novel quantum-mechanical phenomenon known as the quantum shape effect occurs when the geometry of the box potential is modified while keeping all the Lebesgue size parameters, as well as the topology and boundary curvature of the confinement domain, unchanged [51, 52, 53]. The quantum shape effect is achieved through a geometric technique called size-invariant shape transformation [54], illustrated in Fig. 1.
@@
In this study, we investigate the condensation of a non-interacting Bose gas with a finite number of particles confined in a 2D flat-bottom potential between nested square domains with impenetrable boundaries. The angle between the sides of the inner and outer squares defines a size-invariant shape parameter. We demonstrate that quasistatic changes in a shape parameter can induce the BEC. The condensate fraction can be changed and controlled just by shape, without altering the size parameters of the confinement domain, as well as the density and temperature of the gas. Then, we investigate the behavior of Bose-Einstein gas under different dimensionless temperatures and densities to find the optimum condition where the change in condensate fraction due to shape effects is maximized. We calculate the condensate fraction and heat capacity of a Bose gas changing with shape, temperature, and density. Note that our focus is solely on demonstrating that shape alone can induce the condensation of a Bose gas. As long as the overall confinement is sufficiently strong, the effects of geometry can be made dominant over inter-particle interactions [57, 58, 59, 52]. Consequently, we do not employ the Gross-Pitaevskii equation, as the dynamics or other properties of the condensate are beyond the scope of this study.
@@
II Confinement geometry
A non-interacting Bose gas is confined in the region between two nested square domains as shown in Fig. 1 at $\tilde{T}=1$ and $\tilde{n}_{r}=1$, with the potential being infinite inside the inner square and outside the outer square, and zero in between. The outer boundaries of the domain are defined by the side length $L_{o}=L$, while the inner boundaries are set to $L_{i}=0.7L$. The shape control is governed by the parameter $\theta$, which represents the rotation angle of the inner core, allowing us to make size-invariant changes to the domain. Since our investigation focuses explicitly on the BEC transition, we employ a quantum statistical mechanical framework. The confined Bose gas obeys the Bose-Einstein distribution, $f=1/\{\exp[(E-\mu)/(k_{B}T)]-1\}$. There is no explicit time variable, and all changes to the system are performed quasistatically to ensure that thermal equilibrium is maintained at all times. Hence, to determine the energy eigenvalues, we numerically solved the time-independent Schrödinger equation, $\hat{H}\Psi=E\Psi$, for the domain $\Omega$. The Hamiltonian of the system includes explicit dependence on the shape parameter, $\theta$,
@@ II Confinement geometry
To make our results independent of the specific type of Bose gas and any other free parameters, we employed dimensionless values throughout this work. Using these initial values, we calculated the discrete energy eigenvalues for each degree step from $\theta=0$ to $\theta=45$ degrees. We use 5000 of them for each case to ensure that the partition function saturates due to the soft cutoff provided by the Bose function. In doing so, we obtain the energy spectrum for each angular configuration. Subsequently, we converted these results into dimensionless values to generalize our findings and eliminate dependence on the chosen initial parameters.
The dimensionless temperature is defined as $\tilde{T}=T/T_{0}$, where $T_{0}$ is the reference temperature obtained from the ground state energy $E_{0}$ at $\theta=45^{\circ}$ configuration (i.e. minimum $E_{0}$ value), $T_{0}=E_{0}/k_{B}$. Energies are normalized by $k_{B}T$ as, $\tilde{E}=E/(k_{B}T)$, and the dimensionless density is defined as $\tilde{n}_{r}=n/n_{q}(T_{0})=\lambda_{th}^{2}(T_{0})N/A$, where $n=N/A$ is classical surface density, $N$ is the number of particles and $A$ is the area of the confinement domain $\Omega$. Here, $n_{q}(T)=1/\lambda_{th}^{2}(T)$ is the quantum density. The dimensionless density is renormalized using the quantum density at the reference temperature to ensure that the density, and thus the number of particles remain constant for different temperatures. The data and the codes are given in Ref. [61].
III Condensate fraction
Fig. 1 provides qualitative support for the appearance of condensation with increasing $\theta$, indicated by the increased probability density of the ground state. To analyze the BEC transition due to shape effect quantitatively, we determine the condensate fraction,
@@ III Condensate fraction
where $N_{GS}$ is the number of particles in the ground state. $\sum_{E_{GS}}$ represents the sum over the (fourfold) degenerate ground state energy levels, whereas $\sum_{E}$ represents the sums over all energy levels. We calculate chemical potential numerically from the particle number equation, $\sum_{E}f$, as an inverse solution by fixing the particle number and temperature.
@@
[…]
IV Specific heat
Next, we examine the dimensionless specific heat capacity at constant Weyl size parameters (volume, area, periphery, vertices), denoted by $W$ subscript in Eq. (3), under the change of $\theta$ for dimensionless density $\tilde{n}_{r}=1$ and dimensionless temperatures $\tilde{T}=\{0.1,0.5,1,5\}$ as seen in Fig. 3a. The specific heat exhibits distinct and intriguing characteristics that differ significantly from those observed in Bose gases confined in regular flat-bottom geometries or harmonic traps, even in its temperature variation, Fig. 3b.
@@ IV Specific heat
V Conclusion
In conclusion, we have introduced the phenomenon of shape-induced BEC. Size-invariant shape transformations lead to unique behaviors in the condensate fraction and specific heat, offering additional control over the properties of Bose-Einstein condensates. Our findings reveal that the condensate fraction is significantly influenced by the shape parameter. Moreover, we have observed that the shape-dependent specific heat exhibits markedly different behaviors compared to conventional non-interacting Bose gases confined in other potentials.
@@ V Conclusion
VI Acknowledgements
A.A. acknowledges financial support from the Sabanci University Integration Projects Support with project code B.A.CF-24-02877.